# Scale-Dependent Radial Geometry and Metric Mismatch in Wasserstein Propagation for Reverse Diffusion

### 20 Mar 2026 v1

@@ G.4 Worked example: a variance-preserving schedule with Euler/Heun-type defects
$\beta\int_{0}^{s_{0}}\frac{Me^{-\beta u}}{D(u)^{2}}\,\mathop{}\!\mathrm{d}u=\mathcal{M}_{\mathrm{VP}}(s_{0}),$
with the stated formula for $\alpha\neq 1$, and the $\alpha=1$ case obtained by direct integration.
For the discretization terms, the uniform one-step bound gives
$\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q}\sum_{k=0}^{K-1}1\leq C_{\mathrm{sch}}(T-s_{0})h^{q-1},$
and similarly
$\sum_{k=K}^{N-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}\,s_{0}\,h^{q-1}\,\sup_{u\in[0,s_{0}]}e^{\Gamma_{\mathrm{pr}}(u)}.$
Because $\bar{\varepsilon}$ is constant, the early forcing term is
$\int_{s_{0}}^{T}e^{-c_{\mathrm{pr}}(s_{0})(s-s_{0})}\,\beta\,\bar{\varepsilon}\,\mathop{}\!\mathrm{d}s=\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})(T-s_{0})}\Bigr).$
The stated one-switch bound then follows from Proposition 6 and Theorem 1, applied with the proxy inputs. The direct Euclidean comparison is the corresponding specialization of the full-horizon Euclidean bound. ∎

### 30 Mar 2026 v2

@@ G.4 Variance-preserving schedules: explicit routed-versus-direct certificates
$\beta\int_{0}^{s_{0}}\frac{Me^{-\beta u}}{D(u)^{2}}\,\mathop{}\!\mathrm{d}u=\mathcal{M}_{\mathrm{VP}}(s_{0}).$
This gives the stated closed form, with the $\alpha=1$ branch obtained by direct integration.
For (iii), the late-window proxy term is just the split late contribution under the same conservative replacements. On the early window,
$\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q}\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}.$
Since $t_{s}=t_{K}=Kh$,
$t_{s}-t_{k+1}=(K-k-1)h,$
and therefore
$\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}=\sum_{j=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})jh}=\frac{1-e^{-c_{\mathrm{pr}}(s_{0})L}}{1-e^{-c_{\mathrm{pr}}(s_{0})h}}.$
Because $\bar{\varepsilon}$ is constant,
$\int_{s_{0}}^{T}e^{-c_{\mathrm{pr}}(s_{0})(u-s_{0})}\,\beta\bar{\varepsilon}\,\mathop{}\!\mathrm{d}u=\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})L}\Bigr).$
Substituting these estimates into Proposition 6 proves the routed proxy certificate.
For (iv), $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})$ is the direct full-horizon Euclidean certificate obtained by making the same conservative replacements in Proposition 5 and splitting at the same index $K$. If $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})>0$, then for every $r\in[s_{0},T]$,
$\Gamma_{\mathrm{pr}}(r)-\Gamma_{\mathrm{pr}}(s_{0})=\int_{s_{0}}^{r}\mathfrak{b}_{\mathrm{pr}}(u)\,\mathop{}\!\mathrm{d}u\geq\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})\,(r-s_{0}).$
Applying this to $r=T$, $r=T-t_{k+1}$, and inside the forcing integral gives
$e^{\Gamma_{\mathrm{pr}}(T)}\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L},$
$\sum_{k=0}^{K-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\sum_{j=0}^{K-1}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})jh}=e^{\Gamma_{\mathrm{pr}}(s_{0})}\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})h}-1},$
and
$\int_{s_{0}}^{T}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\int_{0}^{L}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})v}\,\mathop{}\!\mathrm{d}v=e^{\Gamma_{\mathrm{pr}}(s_{0})}\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})}.$
This proves the lower bound on $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})-\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})$, and the strict comparison criterion follows immediately. ∎
Remark 13 (Long mismatch windows).
This example makes the advantage of routing completely explicit. If $c_{\mathrm{pr}}(s_{0})L\gg 1$, then the routed early budget has already saturated: the terminal mismatch carries the factor $e^{-c_{\mathrm{pr}}(s_{0})L}$, the pre-switch discretization term is of order $C_{\mathrm{sch}}h^{q-1}/c_{\mathrm{pr}}(s_{0})$, and the forcing term is of order $\beta\bar{\varepsilon}/c_{\mathrm{pr}}(s_{0})$. By contrast, if $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})>0$, then the direct Euclidean certificate transports the same three inputs with positive exponential weights over the same window. So on a genuine mismatch window—radial damping already active, residual Euclidean load still positive—the routed certificate can be strictly sharper, provided the one-time conversion cost $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})$ is not too large.